OpenStudy (angelinaxoxo1):

Solve each proportion for the missing variable. 6). x over 18 = 9 over 27 7). 15 over 25 = 10 over p

OpenStudy (erak):

\[\frac{ x }{ 18 } = \frac{ 9 }{ 27 }\]Basically you would just cross multiply them. That means taking x * 27, and 9 * 18, leaving you with\[27x = 162\]Then solve for x.